DT2DV(1) DT2DV(1) NAME dt2dv−convertaDTLtextrepresentationofaTeXDVIfiletoabinaryDVIfile SYNOPSIS dt2dv[−debug][−group][−si][−so][input-DTL-file][output-DVI-file] Intheabsenceofthe−siand−sooptions,bothfileargumentsarerequiredintheorderinput-DTL-fileoutput-DVI-file.ButalsoseetheOPTIONSsectionbelow.Nodefaultfileexten-sionsaresupplied. DESCRIPTION dt2dvconvertsatextrepresentationofaTEXDVIfile,usuallyproducedbythecompanion dv2dt(1)utility,backtoabinaryDVIfile.DTL(DVITextLanguage)filescanbeedited,with care,andthenrestoredtoDVIformforprocessingbyanyTEXDVIdriverprogram.InDTL files,fontdirectorynamesandfontnamesareprecededbyalengthfield,whichmustbeupdated ifthenamesaremodified. dvitype(1)canalsodisplayatextualrepresentationofDVIfiles,butinsomeimplementationsat least,itcannotbeusedinbatchmode,anditsoutputisnotwell-suitedforconversionbacktoa DVIfile. TheformatofTEXDVIfilesisfullydescribedinDonaldE.Knuth,TEX:TheProgram,Addison-Wesley(1986),ISBN0-201-13437-3,aswellasinthedvitype(1)literateprogramsourcecode. BriefdescriptionsoftheDTLandDVIformatsaregivenindv2dt(1). OPTIONS −debugTurnondetaileddebuggingoutput. −groupExpecteachDTLcommandtobeinparentheses. −siReadallDTLcommandsfromstandardinput. −soWriteallDVIcommandstostandardoutput. SEEALSO dv2dt(1),dvitype(1),tex(1). FILES *.dvibinaryTEXDVIfile. *.dtltextrepresentationofaTEXDVIfileinDVITextLanguageformat. AUTHOR dt2dvanddv2dt(1)werewrittenby GeoffreyTobin DepartmentofElectronicEngineering LaTrobeUniversity Bundoora,Victoria3083 Australia Tel:+6134793736 FAX:+6134793025 Email:<G.Tobin@ee.latrobe.edu.au> Thesemanualpageswerewrittenprimarilyby NelsonH.F.Beebe,Ph.D. CenterforScientificComputing DepartmentofMathematics UniversityofUtah SaltLakeCity,UT84112 Tel:+18015815254 FAX:+18015814148 Email:<beebe@math.utah.edu> Version0.6.0 08March1995 1